OpenStudy (anonymous):

medal for the CORRECT answer! Solve: -4(2x - 1) = -6(x + 2) - 2

OpenStudy (nurali):

-4(2x - 1) = -6(x + 2) - 2 -8x+4=-6x-12-2 -8x+4=-6x-14 -8x+6x=-14-4 -2x=-18 x=9

OpenStudy (anonymous):

Thank you!

OpenStudy (nurali):

My Pleasure.